Question 11 pt Progress: 0/1 item Direct Labor Variances Bellingham Company produces a product that requires 2 standard direct labor hours per unit at a standard hourly rate of $19.00 per hour. If 3,300 units used 6,800 hours at an hourly rate of $18.24 per hour, what is the direct labor (a) rate variance, (b) time variance, and (c) cost variance? Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

$ a. Direct labor rate variance

$ b. Direct labor time variance

$ c. Direct labor cost variance $
Answer #1
Direct labor rate variance = ( Actual rate - Standard rate ) * Actual hours = ( 18.24 - 19.00 ) * 6800 -5168 Favorable
Standard hours for actual production = Actual production * Standard direct labor hour per unit = 3300 * 2 6600
Direct labor time variance = ( Actual hours - Standard hours ) * Standard rate = ( 6800 - 6600 ) * 19.00 3800 Unfavorable
Actual cost of direct labor = Actual hours * Actual rate = 6800 * 18.24 124032
Standard cost of direct labor = Standard hours * Standard rate = 6600 * 19.00 125400
Total direct labor cost variance = Actual cost - Standard cost = 124032 - 125400 -1368 Favorable